OpenStudy (anonymous):

solve the equation sin^2x=3cos^2x

OpenStudy (anonymous):

\[\sin^2(x)=3\cos^2(x)\] On the other hand we know that \[\sin^2(x)+\cos^2(x)=1 \rightarrow \sin^2(x)=1-\cos^2(x)\] Lets put that into the first equation and we get \[1-\cos^2(x)=3\cos^2(x)\] From there on it's fairly simple if you remember the basic angle/length relations with cos/sin/tan functions.